# GS2Pose: Two-stage 6D Object Pose Estimation Guided by Gaussian Splatting

### v2

@@ 1 Introduction
Accurate 6D object pose estimation is a fundamental problem in the field of computer vision, with broad application prospects in technologies such as robot navigation[1, 2] and virtual reality[3, 4]. However, classical pose estimation algorithms[5, 6, 7] lack robustness against environmental interference, such as non-uniform lighting, varying degrees of occlusion, and dynamic blur. Moreover, the lightweight of the algorithm is also demanding in the field of embodied intelligence[8, 9, 10].
@@
With the widespread application of deep learning methods, the robustness of related algorithms[11, 12, 13, 14, 15] against interference has continually improved. Early works[16, 17, 18, 19] have achieved high-precision instance-level pose estimation. However, these models can only handle a specific object after training session and cannot generalize to others. Additionally, they require datasets with precise ground truth pose, which is difficult to obtain in practical applications.
@@
With the emergence of novel pose representation methods[20] such as NOCS, breakthroughs have been made in category-level pose estimation methods[21, 22, 23, 24, 25, 26], achieving notable intra-class generalization. Trained models can perform high-precision pose estimation on objects with similar geometric and color features. However, these methods typically require a substantial amount of CAD models of the same category during training phase, results in huge time expenditure. Additionally, since the 6D pose of the target object is bound to the objects coordinate system under the CAD model, which can lead to issues, such as parameter ambiguity in the estimation results during the inference phase.
@@
In recent years, with the development of large models[27, 28, 29], some research[30, 31, 32] have introduced the concept of pre-training on large datasets into the field of 6D pose estimation. These methods construct large datasets by collecting numerous CAD models of common objects from different categories, enabling effective generalization to unseen objects. They require only the CAD model of the target object during inference, allowing for the artificial setting of strict coordinate relationships without the need for additional training on the target object. However, these models also have drawbacks, such as the inability to generalize to uncommon objects, high computational resource consumption, and their accuracy being heavily dependent on the quality of CAD modeling.
@@
To address the aforementioned shortcomings of these algorithms, we propose a novel pose estimation method that does not require artificial designed CAD models. This method is designed for application scenarios where high-quality CAD models of the target object are unavailable, and only untextured scanned models or SFM point cloud models can be obtained. To achieve lightweight training, accurate reference relationships, and robustness to interference, GS2POSE consists of a two-stage pose estimation approach comprising coarse estimation followed by pose refinement.
@@
The detailed process of GS2Pose is illustrated in in Fig. 1 The 3DGS point cloud model of the object (hereafter referred to as the 3DGS model) is obtained using existing 3DGS reconstruction techniques, with the object coordinate system manually specified. Inspired by the GS-SLAM model[33], we introduce the commonly used reprojection-based pose optimization iterative approach from the SLAM domain[34, 35, 36], also known as Bundle Adjustment (BA). By representing object poses using Lie algebra and integrating it with the differentiable 3DGS rendering pipeline, we developed an approach that utilizes reprojection and backpropagation. It enables an iterative optimization algorithm that can regress both the object pose and the camera pose, referred to as GS-Refiner.
@@
Since this iterative optimization algorithm requires a reasonable initial pose as a starting point for optimization, we also need to design an algorithm that can provide a rough pose estimate using only the segmented object image. Inspired by the NeRF-Pose model[37], we designed a rough pose estimation network called Pose-Unet.RGB images and their corresponding NOCS images are obtained from the camera perspective using 3DGS. These images are subsequently input into a pre-trained coarse pose estimation network (Pose-Unet) for fine-tuning, resulting in a coarse pose estimation for any novel rendering view of the object. .
@@
On the other hand, GS-Refiner leverages the parameter interpretability of the 3DGS model to selectively optimize and refine parameters, such as higher-order spherical harmonic color parameters, transparency, and ellipsoid orientation through backpropagation. This allows the model surface colors to adaptively adjust to environmental factors encountered during actual capture, such as lighting, occlusion, and motion blur.
@@
ii) By employing Lie algebra to modify the differentiable rendering pipeline of 3DGS, a re-projection iterative algorithm called GS-Refiner was developed, enabling the correction of both object poses and camera poses.
@@
iv) Through experiments on datasets such as LineMod, GS2POSE model demonstrated substantial advantages over comparable algorithms, particularly in accuracy, inference speed, and computational resource efficiency.
@@ 2 Related Works
2.1 6D pose estimation of known objects
@@
2.2 6D pose estimation of novel objects
To improve the generalization ability and robustness of 6D pose estimation, some researchers aim to address pose estimation for novel objects. MegaPose[31] network proposed a 6D pose estimator based on a rendering and comparison strategy, which trains the network on a large synthetic dataset. Moreover, GigaPose[30] network proposed a novel solution by leveraging templates to recover out-of-plane rotations, then utilizing patches correspondences to estimate the four remaining pose parameters. Although above foundation methods have strong generalization capabilities, their robustness remains insufficient for specialized devices in industries and medical, such as surgical instruments and precision constructions. We proposed a course-refine 6D pose estimation network. For each new object, coarse estimation network provides an approximate pose by rapid training, followed by precise correction in the refine estimation network.
2.3 3D Gaussian Representation

### v3

@@ 1 Introduction
Accurate 6D object pose estimation is a fundamental problem in the field of computer vision, with broad application prospects in technologies such as robot navigation[1, 2] and virtual reality[3, 4]. However, classical pose estimation algorithms[5, 6, 7] lack robustness against environmental interference, such as non-uniform lighting, varying degrees of occlusion, and dynamic blur. Moreover, the lightweight nature of the algorithm is also demanding in the field of embodied intelligence[8, 9, 10].
@@
With the widespread application of deep learning methods, the robustness of related algorithms[11, 12, 13, 14, 15] against interference has continually improved. Early works[16, 17, 18, 19] have achieved high-precision instance-level pose estimation. However, these models can only handle a specific object after the training session and cannot generalize to others. Additionally, they require datasets with precise ground truth poses, which are difficult to obtain in practical applications.
@@
The emergence of novel pose representation methods [20], such as NOCS, has led to breakthroughs in category-level pose estimation methods[21, 22, 23, 24, 25, 26], achieving notable intra-class generalization. Trained models can perform high-precision pose estimation on objects with similar geometric and color features. However, these methods typically require a substantial number of CAD models of the same category during the training phase, results in huge time expenditure. Additionally, since the 6D pose of the target object is bound to the objects coordinate system under the CAD model, which can lead to issues, such as parameter ambiguity in the estimation results during the inference phase.
@@
In recent years, with the development of large models[27, 28, 29], some research[30, 31, 32] have introduced the concept of pre-training on large datasets into the field of 6D pose estimation. These methods construct large datasets by collecting numerous CAD models of common objects from different categories, enabling effective generalization to unseen objects. They require only the CAD model of the target object during inference, allowing for the artificial setting of strict coordinate relationships without the need for additional training on the target object. However, these models also have drawbacks, such as the inability to generalize to uncommon objects, high consumption of computational resources, and their accuracy being heavily dependent on the quality of CAD modeling.
@@
To address the aforementioned shortcomings of existing algorithms, a novel pose estimation method is proposed that eliminates the need for artificially designed CAD models. This method is designed for application scenarios where high-quality CAD models of the target object are unavailable, and only untextured scanned models or structure-from-motion (SFM) point cloud models can be obtained. To achieve lightweight training, accurate reference relationships, and robustness to interference, GS2Pose consists of a two-stage pose estimation approach comprising coarse estimation followed by pose refinement.
@@
The detailed process of GS2Pose is illustrated in Fig. 1. The 3DGS point cloud model of the object (hereafter referred to as the 3DGS model) is obtained using existing 3DGS reconstruction techniques, with the object coordinate system manually specified. Utilizing insights from the GS-SLAM model[33], the commonly used reprojection-based pose optimization iterative approach from the SLAM domain is introduced [34, 35, 36], also known as Bundle Adjustment (BA). By representing object poses using Lie algebra and integrating this representation with the differentiable 3DGS rendering pipeline, an approach is implemented that utilizes reprojection and backpropagation. This enables an iterative optimization algorithm that can regress both the object pose and the camera pose, referred to as GS-Refiner.
@@
Since the iterative optimization algorithm requires a reasonable initial pose as a starting point, it is necessary to design an algorithm that can provide a rough pose estimate based solely on the segmented object image. Inspired by the NeRF-Pose model[37], a rough pose estimation network named Pose-Unet was developed. RGB images and their corresponding NOCS images are obtained from the camera perspective using 3DGS. These images are subsequently input into a pre-trained coarse pose estimation network (Pose-Unet) for fine-tuning, resulting in a coarse pose estimation for any novel rendering view of the object.
@@
On the other hand, GS-Refiner leverages the parameter interpretability of the 3DGS model to selectively optimize and refine parameters, such as higher-order spherical harmonic color parameters, transparency, and ellipsoid orientation through backpropagation. This allows the surface colors to adaptively adjust to environmental factors encountered during actual capture, such as lighting, occlusion, and motion blur.
@@
ii) By employing Lie algebra to modify the differentiable rendering pipeline of 3DGS, a reprojection iterative algorithm called GS-Refiner has been developed developed, enabling the correction of both object poses and camera poses.
@@
iv) Through experiments on datasets such as LineMod, the GS2Pose model demonstrated substantial advantages over comparable algorithms, particularly in terms of accuracy, inference speed, and computational resource efficiency.
@@ 2 Related Works
2.1 6D pose estimation of seen objects
@@
2.2 6D pose estimation of unseen objects
To improve the generalization ability and robustness of 6D pose estimation with CAD models, some researchers aim to address pose estimation for novel objects. MegaPose[31] network proposed a 6D pose estimator based on a rendering and comparison strategy, which trains the network on a large synthetic dataset. Moreover, GigaPose[30] network proposed a novel solution by leveraging templates to recover out-of-plane rotations, then utilizing patches correspondences to estimate the four remaining pose parameters. Although above foundation methods have strong generalization capabilities, their robustness remains insufficient for specialized devices in industries and medical, such as surgical instruments and precision constructions. We proposed a course-refine 6D pose estimation network. For each new object, coarse estimation network provides an approximate pose by rapid training, followed by precise correction in the refine estimation network.
2.3 6D pose estimation with 3D reconstruction model